OpenStudy (anonymous):

how to write; 2x+3y=12 in slope intercept form

OpenStudy (anonymous):

Slope intercept form is y=mx+b Try to keep the y term to the left side and get the x-term on the right. What do you get? Then isolate the y by dividing the answer by 3

OpenStudy (anonymous):

subtract 2x from each side... then divide each side by 3

OpenStudy (anonymous):

2x+3y=12 -2x -2x ______________ 3y= -2x+12 3y/3 =-2x/ 3 + 12/3 y=-2/3x=4 Need me to explain further or any more help?

OpenStudy (anonymous):

ThailynnIsMe if you have to subtract 2x from both sides how come you put 3y= -2x=12 shouldnt it be 12-2?

OpenStudy (anonymous):

Sorry the answer is y=-2/3x+4 It was a typo! Sorry :)

OpenStudy (anonymous):

Do you have it now?

OpenStudy (anonymous):

yeah, kinda

OpenStudy (anonymous):

I'll explain with words if you want and show you even slower and point it out if need be

OpenStudy (anonymous):

okay this is what i got: 2x+3y+12 -2x -2x _____________ 3y+10 3/3 10/3 y= 3.33 repeating. am i way off?

OpenStudy (anonymous):

Are you trying to subtract 2x from both sides because it does not work like that. I'm going to show you again step by step. Hold on

OpenStudy (anonymous):

*2x+3y=12

OpenStudy (anonymous):

Alright hold on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Almost done

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

Got it, sorry my computer was acting crazy

OpenStudy (anonymous):

There you go, just open the attachment! :)